OpenStudy (pulsified333):

involving the sum of the numbers showing on two fair dice. (1) What is the probability that exactly one die shows a 6 given that the sum of the numbers is 9? (2) What is the probability that the sum of the numbers is 9 given that exactly one die shows a 6? (3) What is the probability that the sum of the numbers is 9 given that at least one die shows a 6?

OpenStudy (pulsified333):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

which one are you stuck on? how far did you get?

OpenStudy (pulsified333):

i haven't gotten any yet. lets start with #1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it tells us in #1, that `given that the sum of the numbers is 9` so we know, for a fact, that the two dice add to 9

jimthompson5910 (jim_thompson5910):

what are the ways to add to 9 with 2 dice?

OpenStudy (pulsified333):

4&5

jimthompson5910 (jim_thompson5910):

yep, what else? can you list ALL of the ways to add to 9?

OpenStudy (pulsified333):

3&6, 4&5

jimthompson5910 (jim_thompson5910):

so put that all together we have 3+6 4+5 5+4 6+3 there are only 4 ways to do this

jimthompson5910 (jim_thompson5910):

of these 4 ways, how many have exactly one "6" in them?

jimthompson5910 (jim_thompson5910):

you sure?

OpenStudy (pulsified333):

2?

jimthompson5910 (jim_thompson5910):

yeah 2 3+6 and 6+3

jimthompson5910 (jim_thompson5910):

it would be a fraction Probability of exactly one 6, given sum of 9 = (# of sums with 6 in them)/(# of ways to add to 9) = 2/4 = 1/2

jimthompson5910 (jim_thompson5910):

` given that exactly one die shows a 6` so we know die A is 6 or die B is 6 (both cannot be 6 at the same time)

jimthompson5910 (jim_thompson5910):

if die A is 6, then what are the possibilities for die B?

OpenStudy (pulsified333):

5 and 4?

jimthompson5910 (jim_thompson5910):

B could also be 1,2,3 basically it could be anything but 6

jimthompson5910 (jim_thompson5910):

if die A is 6, then die B could be 1,2,3,4,5 similarly if die B is 6, then die A could be 1,2,3,4,5

OpenStudy (pulsified333):

but doesn't the sum have to be 9?

jimthompson5910 (jim_thompson5910):

no, the given here is that one die is 6. That's it

jimthompson5910 (jim_thompson5910):

this dice chart may help

OpenStudy (pulsified333):

ok so die B could be 1,2,3,4,5

jimthompson5910 (jim_thompson5910):

how many possible outcomes are there if exactly one die is 6?

OpenStudy (pulsified333):

but it says What is the probability that the sum of the numbers is 9

jimthompson5910 (jim_thompson5910):

we'll get there

jimthompson5910 (jim_thompson5910):

in the chart I posted, mark the row that has 6 and the column that has 6. But don't mark the cell that has both 6's. How many cells did you highlight?

OpenStudy (pulsified333):

5?

jimthompson5910 (jim_thompson5910):

you should count 5 along the '6' row and 5 along the '6' column so 10 in total

jimthompson5910 (jim_thompson5910):

how many of those 10, have a sum of 9?

OpenStudy (pulsified333):

2?

jimthompson5910 (jim_thompson5910):

yes, 6+3 and 3+6 we have 2 sums that have a sum of 9 out of 10 outcomes so 2/10 = 1/5 is the answer to #2

OpenStudy (pulsified333):

oh ok, thanks

jimthompson5910 (jim_thompson5910):

and for #3, it's almost identical to #2 BUT the key word in #3 is `at least` so it's possible to have both dice be 6

OpenStudy (pulsified333):

6&3 3&6

OpenStudy (pulsified333):

so is it 1/5? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

we have 2 ways to add to 9 (6+3 or 3+6) however, instead of 10 outcomes, we actually have 11. Again, it's possible to have both dice be 6 so it's actually 2/11

OpenStudy (pulsified333):

thank you :)

jimthompson5910 (jim_thompson5910):

no problem